OpenStudy (anonymous):

A 20 foot ladder is placed against a building to reach a window that is 16 feet above the ground. How many feet away from the building is the bottom of the ladder? A. 3 B. 4 C. 12 D. 15

OpenStudy (anonymous):

Use the Pythagorean theorem. \(a^2 + b^2 = c^2\) 

OpenStudy (anonymous):

i got 4 which is b? correct???

OpenStudy (anonymous):

or no?

OpenStudy (anonymous):

No. You have to square the sides before you add them. \[a^2 + 16^2 = 20^2\]

OpenStudy (anonymous):

okay got it thanks